# Global Convergence for Average Reward Constrained MDPs with Primal-Dual Actor Critic Algorithm

### v1

@@ Abstract
This paper investigates infinite-horizon average reward Constrained Markov Decision Processes (CMDPs) with general parametrization. We propose a Primal-Dual Natural Actor-Critic algorithm that adeptly manages constraints while ensuring a high convergence rate. In particular, our algorithm achieves global convergence and constraint violation rates of $\tilde{\mathcal{O}}(1/\sqrt{T})$ over a horizon of length $T$ when the mixing time, $\tau_{\mathrm{mix}}$, is known to the learner. In absence of knowledge of $\tau_{\mathrm{mix}}$, the achievable rates change to $\tilde{\mathcal{O}}(1/T^{0.5-\epsilon})$ provided that $T\geq\tilde{\operatorname*{\mathcal{O}}}\left(\tau_{\mathrm{mix}}^{2/\epsilon}\right)$. Our results match the theoretical lower bound for Markov Decision Processes and establish a new benchmark in the theoretical exploration of average reward CMDPs.
@@ 1 Introduction
Reinforcement Learning (RL) is a paradigm where an agent learns to maximize its average reward in an unknown environment through repeated interactions. RL finds applications in diverse domains such as transportation, communication networks, robotics, and epidemic control (yang2020deep, ; al2019deeppool, ; ling2023cooperating, ). Among various RL settings, the infinite horizon average reward setup is of particular significance for modeling long-term objectives in practical scenarios. This setting is critical as it aligns with real-world applications requiring persistent and consistent performance over time.
@@
Finding optimal policies for average reward CMDPs is challenging, especially if the environment is unknown. The efficiency of a solution to the CMDP is measured by the rate at which the average global optimality error and the constraint violation diminish as a function of the length of the horizon $T$. Although many existing works focus on the tabular policies agarwal2022concave ; chen2022learning , these solutions do not apply to real-life scenarios where the state space is large or infinite.
@@
UC-CURL and PS-CURL (agarwal2022concave, ) $\tilde{\mathcal{O}}\bigg{(}1/\sqrt{T}\bigg{)}$ 00 Yes No Tabular
@@
Triple-QA (wei2022provably, ) $\tilde{\mathcal{O}}\bigg{(}1/T^{1/6}\bigg{)}$ 00 Yes Yes Tabular
@@
Table 1: This table summarizes the different model-based and model-free state-of-the-art algorithms available in the literature for average reward CMDPs and their results for global convergence rate and average constraint violation.
@@
Our algorithm has a nested loop structure where the outer loop of length $K$ updates the primal-dual parameters, and the inner loops of length $H$ run the natural policy gradient (NPG) and optimal critic parameter finding subroutines. We show that, to achieve the optimal rates, the parameter $H$ should be nearly constant while $K$ should be $\tilde{\Theta}(T)$. This adjustment makes our algorithm resemble a single-timescale algorithm. In contrast, for unconstrained MDPs, $K$ and $H$ are typically set to $\Theta(\sqrt{T})$ ganesh2024accelerated . Achieving $\mathcal{O}(1/\sqrt{T})$ convergence in our setting requires ensuring that the bias remains sufficiently small despite a significantly low value of $H$. Achieving this challenging goal is made possible due to the use of Multi-level Monte Carlo (MLMC)-based estimates in the inner loop subroutines and a sharper analysis of these updates. One benefit of using MLMC-based estimates is that it allows us to design an algorithm that achieves near-optimal global convergence and constraint violation rates without knowledge of the mixing time. We would like to emphasize that our work is the first in the general parameterized CMDP literature to achieve this feat.
@@ 2 Formulation
The mixing time of a CMDP $\mathcal{M}$ with respect to a policy parameter $\theta$ is defined as, $\tau_{\mathrm{mix}}^{\theta}\coloneqq\min\left\{t\geq 1\bigg{|}\|(P^{\pi_{\theta}})^{t}(s,\cdot)-d^{\pi_{\theta}}\|\leq\dfrac{1}{4},\forall s\in\mathcal{S}\right\}$. We also define $\tau_{\mathrm{mix}}\coloneqq\sup_{\theta\in\Theta}\tau^{\theta}_{\mathrm{mix}}$ as the the overall mixing time. In this paper, $\tau_{\mathrm{mix}}$ is finite due to ergodicity.
@@
$\nabla_{\theta}J_{g}(\theta)\!=\!\mathbb{E}_{s\sim d^{\pi_{\theta}},a\sim\pi_{\theta}(\cdot|s)}\!\bigg{[}\!A_{g}^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\!\bigg{]}$ (7)
@@ 3 Algorithm
[…]
@@
where $\alpha$ and $\beta$ denote primal and dual learning rates respectively, and $\delta$ indicates the Slater parameter introduced in Assumption 2.1. Moreover, for any $\Lambda\subset\mathbb{R}$, $\mathcal{P}_{\Lambda}$ indicates the projection onto $\Lambda$. Since $\nabla_{\theta}\mathcal{L}(\theta_{k},\lambda_{k})$, $F(\theta_{k})$, and $J_{c}(\theta_{k})$ are not exactly computable due to lack of knowledge of the exact transition kernel, $P$, and thus, that of the occupation measure, $\nu^{\pi_{\theta_{k}}}$, in most RL scenarios, we use the following approximate updates.
@@ 3.1.1 Critic Estimation
The second task is to estimate the value function $V_{g}^{\pi_{\theta_{k}}}$. To facilitate this objective, it is assumed that, $\forall g\in\{r,c\}$, the value function $V_{g}^{\pi_{\theta_{k}}}(\cdot)$ is well-approximated by a linear critic function $\hat{V}_{g}(\zeta_{g}^{\theta_{k}},\cdot)\coloneqq\langle\phi_{g}(\cdot),\zeta_{g}^{\theta_{k}}\rangle$ where $\phi_{g}:\mathcal{S}\rightarrow\mathbb{R}^{m}$ is a feaure mapping with the property that $\|\phi_{g}(s)\|\leq 1$, $\forall s\in\mathcal{S}$, and $\zeta^{\theta_{k}}_{g}\in\mathbb{R}^{m}$ is a solution of the following optimization.
@@
=$\overset{(a)}{=}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu_{g}^{\pi_{\theta_{k}}}(s,a)\left(\zeta^{\top}\phi_{g}(s)-g(s,a)+J_{g}(\theta_{k})-\mathbb{E}_{s^{\prime}\sim P(s,a)}[V_{g}^{\pi_{\theta_{k}}}(s^{\prime})]\right)\phi_{g}(s)$ (16)
@@
where $c_{\gamma}$ is a constant, $\gamma_{\xi}$ defines a learning parameter, and $h\in\{0,\cdots,H-1\}$. Moreover, the terms ^$\hat{\nabla}_{\eta}R_{g}(\theta_{k},\eta_{g,h}^{k})$ and ^$\hat{\nabla}_{\zeta}E_{g}(\theta_{k},\xi_{g,h}^{k})$ indicate estimates of $\nabla_{\eta}R_{g}(\theta_{k},\eta_{g,h}^{k})$ and $\nabla_{\zeta}E_{g}(\theta_{k},\zeta_{g,h}^{k})$ respectively where $\xi_{g,h}^{k}\triangleq[\eta_{g,h}^{k},(\zeta_{g,h}^{k})^{\top}]^{\top}$. Because the expression of $\nabla_{\zeta}E_{g}(\theta_{k},\zeta_{g,h}^{k})$ comprises $\zeta_{g,h}^{k}$ and $J_{g}(\theta_{k})$ (refer (16)), its estimate is a function of $\eta_{g,h}^{k}$ and $\zeta_{g,h}^{k}$. At the end of inner loop iterations, we obtain $\eta^{k}_{g,H}$ and $\zeta_{g,H}^{k}$ which are estimates of $J_{g}(\theta_{k})$ and $\zeta_{g}^{\theta_{k}}$ respectively. It remains to see how the gradient estimates used in (17) can be obtained. Note that (17) can be compactly written as
@@
where $T_{\max}$ is a constant, $\mathbf{v}^{j}_{g,kh}=2^{-j}\sum_{t=0}^{2^{j}-1}\mathbf{v}_{g}(\theta_{k},\xi_{g,h}^{k};z_{kh}^{t})$, $j\in\{0,Q_{h}^{k}-1,Q_{h}^{k}\}$. Note that the maximum number of state transition samples utilized in the MLMC estimate is $T_{\max}$. Moreover, it can be demonstrated that the samples used on average is $\tilde{O}(\log T_{\max})$. The advantage of the MLMC estimator is that it achieves the same bias as averaging $T_{\max}$ samples, but requires only $\tilde{\operatorname*{\mathcal{O}}}(\log T_{\max})$ samples. In addition, since drawing from a geometric distribution does not require knowledge of the mixing time, we can eliminate the mixing time knowledge assumption used in previous works bai2023regret ; ganesh2024variance ; bai2024learning . Furthermore, these previous works utilizes policy gradient methods and require saving the trajectories of length $H$ for gradient estimations, while our approach does not. Therefore, our algorithm reduces the memory complexity by a factor of $H$, which is significant since the choice of $H$ in these works scales with mixing and hitting times of the MDP.
@@ 3.1.2 Natural Policy Gradient (NPG) Estimator
where ^$\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k})$ is an MLMC-based estimate of $\nabla_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k})$ where $f_{g}$ is given in (9). To obtain this estimate, a $\pi_{\theta_{k}}$-induced trajectory $\mathcal{T}_{kh}=\{(s_{kh}^{t},a_{kh}^{t},s_{kh}^{t+1})\}_{t=0}^{l_{kh}-1}$ of length $l_{kh}=2^{P_{h}^{k}}$ is considered where $P_{h}^{k}\sim\mathrm{Geom}(1/2)$. For a certain transition $z_{kh}^{j}=(s_{kh}^{j},a_{kh}^{j},s_{kh}^{j+1})$, define the following estimate.
^$\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k};z_{kh}^{j})=\hat{F}(\theta_{k};z_{kh}^{j})\omega_{g,h}^{k}-\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z_{kh}^{j})~{}\text{where}~{}$ (22)
@@
^$\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z_{kh}^{j})=\hat{A}_{g}^{\pi_{\theta_{k}}}(\xi_{g}^{k};z_{kh}^{j})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{kh}^{j}|s_{kh}^{j})$
=$\hskip 65.44142pt\overset{(a)}{=}\left[g(s_{kh}^{j},a_{kh}^{j})-\eta_{g}^{k}+\zeta_{g}^{k}\left(\phi_{g}(s_{kh}^{j+1})-\phi_{g}(s_{kh}^{j})\right)\right]\nabla_{\theta}\log\pi_{\theta_{k}}(a_{kh}^{j}|s_{kh}^{j})$
@@
^$\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k})$ $=\mathbf{u}_{g,kh}^{0}+\begin{cases}2^{P_{h}^{k}}(\mathbf{u}_{g,kh}^{P_{h}^{k}}-\mathbf{u}_{g,kh}^{P_{h}^{k}-1}),&\text{if }2^{P_{h}^{k}}\leq T_{\max}\\ 0,&\text{otherwise}\end{cases}$ (23)
@@ 4 Global Convergence Analysis
$\begin{split}&(a)\|\nabla_{\theta}\log\pi_{\theta}(a|s)\|\leq G_{1}\quad\\ &(b)\|\nabla_{\theta}\log\pi_{\theta}(a|s)-\nabla_{\theta}\log\pi_{\theta_{2}}(a|s)\|\leq G_{2}\|\theta_{1}-\theta_{2}\|.\end{split}$
@@
where $(a)$ follows from Assumption 4.5 and the definition that $\omega_{k}^{*}=F(\theta_{k})^{-1}\nabla_{\theta}\mathcal{L}(\theta_{k},\lambda_{k})$. We can obtain a global convergence bound by bounding the terms $\mathbb{E}\|\omega_{k}-\omega^{*}_{k}\|^{2}$, $\mathbb{E}\|(\mathbb{E}\left[\omega_{k}|\theta_{k}\right]-\omega^{*}_{k})\|$ and $\mathbb{E}\|\nabla_{\theta}\mathcal{L}(\theta_{k},\lambda_{k})\|^{2}$. The first two terms are the variance and bias of the NPG estimator, $\omega_{k}$, and the third term indicates the local convergence rate. Further, $\mathbb{E}\|\nabla_{\theta}\mathcal{L}(\theta_{k},\lambda_{k})\|^{2}$ can be upper bounded by a constant (Lemma G.2 in the appendix). We now provide the convergence result for the actor and critic parameters. For brevity, we use $\lesssim$ to denote $\leq\tilde{\operatorname*{\mathcal{O}}}(\cdot)$.

### v2

@@ Abstract
This paper investigates infinite-horizon average reward Constrained Markov Decision Processes (CMDPs) under general parametrized policies with smooth and bounded policy gradients. We propose a Primal-Dual Natural Actor-Critic algorithm that adeptly manages constraints while ensuring a high convergence rate. In particular, our algorithm achieves global convergence and constraint violation rates of $\tilde{\mathcal{O}}(1/\sqrt{T})$ over a horizon of length $T$ when the mixing time, $\tau_{\mathrm{mix}}$, is known to the learner. In absence of knowledge of $\tau_{\mathrm{mix}}$, the achievable rates change to $\tilde{\mathcal{O}}(1/T^{0.5-\epsilon})$ provided that $T\geq\tilde{\operatorname*{\mathcal{O}}}\left(\tau_{\mathrm{mix}}^{2/\epsilon}\right)$. Our results match the theoretical lower bound for Markov Decision Processes and establish a new benchmark in the theoretical exploration of average reward CMDPs.
@@ 1 Introduction
Reinforcement Learning (RL) is a paradigm where an agent learns to maximize its average reward in an unknown environment through repeated interactions. RL finds applications in diverse domains such as transportation, communication networks, robotics, and epidemic control (al2019deeppool, ; haliem2021distributed, ; panju2021queueing, ; chen2023option, ; ling2023cooperating, ). Among various RL settings, the infinite horizon average reward setup is of particular significance for modeling long-term objectives in practical scenarios. This setting is critical as it aligns with real-world applications requiring persistent and consistent performance over time.
@@
Finding optimal policies for average reward CMDPs is challenging, especially if the environment is unknown. The efficiency of a solution to the CMDP is measured by the rate at which the average global optimality error and the constraint violation diminish as a function of the length of the horizon $T$. Although many existing works focus on the tabular policies gattami2021reinforcement ; agarwal2022regret ; agarwal2022concave ; chen2022learning , these solutions do not apply to real-life scenarios where the state space is large or infinite.
@@
UC-CURL and PS-CURL (agarwal2022concave, ) $\tilde{\mathcal{O}}\bigg(1/\sqrt{T}\bigg)$ 0 Yes No Tabular
@@
Triple-QA (wei2022provably, ) $\tilde{\mathcal{O}}\bigg(1/T^{1/6}\bigg)$ 0 Yes Yes Tabular
@@
Table 1: This table summarizes the different model-based and model-free state-of-the-art algorithms available in the literature for average reward CMDPs and their results for global convergence rate and average constraint violation. The bounds of global convergence describe convergence to the best performance permitted by the chosen features and policy class; the residual approximation floor is fixed (independent of $T$), and the listed rates govern how fast the statistical gap decays toward that floor. General parameterization refers to parameterizations whose policy score $\nabla_{\theta}\log\pi_{\theta}(a|s)$ is uniformly bounded and Lipschitz in $\theta$, together with a Fisher non-degeneracy condition.
@@
Our algorithm has a nested loop structure where the outer loop of length $K$ updates the primal-dual parameters, and the inner loops of length $H$ run the natural policy gradient (NPG) and optimal critic parameter finding subroutines. We show that, to achieve the optimal rates, the parameter $H$ should be nearly constant while $K$ should be $\tilde{\Theta}(T)$. This adjustment makes our algorithm resemble a single-timescale algorithm. In contrast, for unconstrained MDPs, $K$ and $H$ are typically set to $\Theta(\sqrt{T})$ ganesh2024accelerated . Achieving $\mathcal{O}(1/\sqrt{T})$ convergence in our setting requires ensuring that the bias remains sufficiently small despite a significantly low value of $H$. Achieving this challenging goal is made possible due to the use of Multi-level Monte Carlo (MLMC)-based estimates in the inner loop subroutines and a sharper analysis of these updates. One benefit of using MLMC-based estimates is that it allows us to design an algorithm that achieves near-optimal global convergence and constraint violation rates without knowledge of the mixing time. Related unconstrained average-reward results have also used MLMC to avoid mixing-time oracles (suttle2023beyond, ; patel2024towards, ; ganesh2024accelerated, ); here we apply this idea to CMDPs via a primal–dual natural actor–critic approach and provide global-rate and average-violation guarantees. We would like to emphasize that our work is the first in the general parameterized CMDP literature to achieve this feat.
@@ 2 Formulation
[…]
@@
$\nabla_{\theta}J_{g}(\theta)\!=\!\mathbb{E}_{\begin{subarray}{c}s\sim d^{\pi_{\theta}}\\ a\sim\pi_{\theta}(\cdot\mid s)\end{subarray}}\!\bigg[\!A_{g}^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\!\bigg]$ (7)
@@ 3 Algorithm
We solve (3) via a primal-dual algorithm based on the following saddle point optimization.
@@
where $\alpha$ and $\beta$ denote primal and dual learning rates respectively, and $\delta$ indicates the Slater parameter introduced in Assumption 2.1. Moreover, for any $\Lambda\subset\mathbb{R}$, $\mathcal{P}_{\Lambda}$ indicates the projection onto $\Lambda$. Since $\nabla_{\theta}\mathcal{L}(\theta_{k},\lambda_{k})$, $F(\theta_{k})$, and $J_{c}(\theta_{k})$ are not exactly computable due to lack of knowledge of the exact transition kernel, $P$, and thus, that of the occupancy measure, $\nu^{\pi_{\theta_{k}}}$, in most RL scenarios, we use the following approximate updates.
@@ 3.1.1 Critic Estimation
The second task is to estimate the value function $V_{g}^{\pi_{\theta_{k}}}$. To facilitate this objective, it is assumed that, $\forall g\in\{r,c\}$, the value function $V_{g}^{\pi_{\theta_{k}}}(\cdot)$ is well-approximated by a linear critic function $\hat{V}_{g}(\zeta_{g}^{\theta_{k}},\cdot)\coloneqq\langle\phi_{g}(\cdot),\zeta_{g}^{\theta_{k}}\rangle$ where $\phi_{g}:\mathcal{S}\rightarrow\mathbb{R}^{m}$ is a feature mapping with the property that $\|\phi_{g}(s)\|\leq 1$, $\forall s\in\mathcal{S}$, and $\zeta^{\theta_{k}}_{g}\in\mathbb{R}^{m}$ is a solution of the following optimization.
@@
$\hskip-42.67912pt\overset{(a)}{=}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu_{g}^{\pi_{\theta_{k}}}(s,a)\left(\zeta^{\top}\phi_{g}(s)-g(s,a)+J_{g}(\theta_{k})-\mathbb{E}_{s^{\prime}\sim P(s,a)}[V_{g}^{\pi_{\theta_{k}}}(s^{\prime})]\right)\phi_{g}(s)$ (16)
@@
where $c_{\gamma}$ is a constant, $\gamma_{\xi}$ defines a learning parameter, and $h\in\{0,\cdots,H-1\}$. Moreover, the terms $\hat{\nabla}_{\eta}R_{g}(\theta_{k},\eta_{g,h}^{k})$ and $\hat{\nabla}_{\zeta}E_{g}(\theta_{k},\xi_{g,h}^{k})$ indicate estimates of $\nabla_{\eta}R_{g}(\theta_{k},\eta_{g,h}^{k})$ and $\nabla_{\zeta}E_{g}(\theta_{k},\zeta_{g,h}^{k})$ respectively where $\xi_{g,h}^{k}\triangleq[\eta_{g,h}^{k},(\zeta_{g,h}^{k})^{\top}]^{\top}$. Because the expression of $\nabla_{\zeta}E_{g}(\theta_{k},\zeta_{g,h}^{k})$ comprises $\zeta_{g,h}^{k}$ and $J_{g}(\theta_{k})$ (refer (16)), its estimate is a function of $\eta_{g,h}^{k}$ and $\zeta_{g,h}^{k}$. At the end of inner loop iterations, we obtain $\eta^{k}_{g,H}$ and $\zeta_{g,H}^{k}$ which are estimates of $J_{g}(\theta_{k})$ and $\zeta_{g}^{\theta_{k}}$ respectively. It remains to see how the gradient estimates used in (17) can be obtained. Note that (17) can be compactly written as
@@
where $T_{\max}$ is a constant, $\mathbf{v}^{j}_{g,kh}=2^{-j}\sum_{t=0}^{2^{j}-1}\mathbf{v}_{g}(\theta_{k},\xi_{g,h}^{k};z_{kh}^{t})$, $j\in\{0,Q_{h}^{k}-1,Q_{h}^{k}\}$. Note that the maximum number of state transition samples utilized in the MLMC estimate is $T_{\max}$. Moreover, it can be demonstrated that the samples used on average is $\tilde{O}(\log T_{\max})$. The advantage of the MLMC estimator is that it achieves the same bias as averaging $T_{\max}$ samples, but requires only $\tilde{\operatorname*{\mathcal{O}}}(\log T_{\max})$ samples. In addition, since drawing from a geometric distribution does not require knowledge of the mixing time, we can eliminate the mixing time knowledge assumption used in previous works bai2023regret ; ganesh2024variance ; bai2024learning . Furthermore, these previous works utilize policy gradient methods and require saving the trajectories of length $H$ for gradient estimations, while our approach does not. Therefore, our algorithm reduces the memory complexity by a factor of $H$, which is significant since the choice of $H$ in these works scales with mixing and hitting times of the MDP.
@@ 3.1.2 Natural Policy Gradient (NPG) Estimator
where $\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k})$ is an MLMC-based estimate of $\nabla_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k})$ where $f_{g}$ is given in (9). To obtain this estimate, a $\pi_{\theta_{k}}$-induced trajectory $\mathcal{T}_{kh}=\{(s_{kh}^{t},a_{kh}^{t},s_{kh}^{t+1})\}_{t=0}^{l_{kh}-1}$ of length $l_{kh}=2^{P_{h}^{k}}$ is considered where $P_{h}^{k}\sim\mathrm{Geom}(1/2)$. For a certain transition $z_{kh}^{j}=(s_{kh}^{j},a_{kh}^{j},s_{kh}^{j+1})$, define the following estimate.
$\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k};z_{kh}^{j})=\hat{F}(\theta_{k};z_{kh}^{j})\omega_{g,h}^{k}-\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z_{kh}^{j})~\text{where}$~ (22)
@@
$\hat{\nabla}_{\theta}J_{g}(\theta_{k},\xi_{g}^{k};z_{kh}^{j})=\hat{A}_{g}^{\pi_{\theta_{k}}}(\xi_{g}^{k};z_{kh}^{j})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{kh}^{j}|s_{kh}^{j})$
$\hskip 65.44142pt\overset{(a)}{=}\left[g(s_{kh}^{j},a_{kh}^{j})-\eta_{g}^{k}+\zeta_{g}^{k}\left(\phi_{g}(s_{kh}^{j+1})-\phi_{g}(s_{kh}^{j})\right)\right]\nabla_{\theta}\log\pi_{\theta_{k}}(a_{kh}^{j}|s_{kh}^{j})$
@@
$\hat{\nabla}_{\omega}f_{g}(\theta_{k},\omega_{g,h}^{k},\xi_{g}^{k})$ $=\mathbf{u}_{g,kh}^{0}+\begin{cases}2^{P_{h}^{k}}(\mathbf{u}_{g,kh}^{P_{h}^{k}}-\mathbf{u}_{g,kh}^{P_{h}^{k}-1}),&\text{if }2^{P_{h}^{k}}\leq T_{\max}\\ 0,&\text{otherwise}\end{cases}$ (23)
@@ 4 Global Convergence Analysis
$\begin{split}&(a)\|\nabla_{\theta}\log\pi_{\theta}(a|s)\|\leq G_{1}\quad\\ &(b)\|\nabla_{\theta}\log\pi_{\theta_{1}}(a|s)-\nabla_{\theta}\log\pi_{\theta_{2}}(a|s)\|\leq G_{2}\|\theta_{1}-\theta_{2}\|.\end{split}$
@@
[…]
@@
In the setting of Theorem 4.10, for small $\epsilon$, the horizon requirement becomes $T\gtrsim\tilde{O}\left(\tau_{\mathrm{mix}}^{2/\varepsilon}\right)$, which can be large for slowly mixing problems. In such a scenario, one can switch to the parameter setting of Theorem 4.9, which does not impose any such restriction on $T$. However, since this setup requires the knowledge of $\tau_{\rm mix}$, one can treat $\tau_{\rm mix}$ as one of the unknown hyperparameters of the algorithm, and fine-tune it during the training phase. This is in line with other RL algorithms in the literature that also require fine-tuning of several unknown hyperparameters, such as the Lipschitz constant or hitting time (bai2023regret, ). Despite these practical solutions, we acknowledge that a systematic theoretical investigation is needed to improve the requirement of the horizon length, $T$ (in the absence of knowledge of $\tau_{\rm mix}$), which is left as one of the future works.